If I am at the gas station filling my tank with gas. The function C(g) represents the cost C of filling up the gas tank with g gallons. Given the equation, what does the number 3.03 represent? What would C(2) or C(9) be?
If I graph C(g), what would be an appropriate domain and range be?

Answer by stanbon(75887)


C(g)=3.03(g)
If I am at the gas station filling my tank with gas. The function C(g) represents the cost C of filling up the gas tank with g gallons. Given the equation, what does the number 3.03 represent? What would C(2) or C(9) be?
---------------------
3.03 is the slope of the equation and is the increased
price of the cost (C) for each additional gallon of gas.
----------
C(2) = 3.03*2 = $6.06
-------
C(9) = 3.03*9 = $27.27
--------
Cheers,
Stan H.